Convert the expression into postfix notation and then verify if it is correct or not,direct answers for both 5conversion are not allowed for verification take dummy values for eachvariable.


a+b*(c/d-e)%(f+g*h)-i


Note:

1)No code is required.


1
Expert's answer
2021-10-05T06:46:52-0400
To compile  postfix notation
Let n is the output string, st is the stack of operations
We read the expression a+b*(c/d-e)%(f+g*h)-i
read a add a to  n= a
read + add + to st st= +
read b add b to  n= ab
read * add * to st st= *+
read ( add ( to st st= (*+
read с add с  to  n= abс
read / add / to st st= /(*+
read d add d  to n   n= abсd
read  - execution priority - less / top stack item
add top stack item to n   n= abсd/
add - to st st= -(*+
read e add e  to n   n= abсd/e
read ) 
We attach the stack to the output line until we get an open parenthesis.
add top stack item to n   n= abсd/e-  st= (*+
top stack item ( delete ( st= *+
read  % execution priority % less * top stack item
add top stack item to n   n= abсd/e-*
add % to st st= %+
read ( add ( to st st= ( %+
read f add f to  n= abсd/e-*f
read + add + to st st= +( %+
read g add g to  n= abсd/e-*fg
read * add * to st st= *+( %+
read h addg to  n= abсd/e-*fgh
read ) 
We attach the stack to the output line until we get an open parenthesis.
n =  abсd/e-*fgh*+
st= %+
read  - execution priority - less % top stack item
add top stack item to n   n =  abсd/e-*fgh*+%
st = + 
top stack item + same priority with -
add top stack item to n   n =  abсd/e-*fgh*+%+
st empty
add - to st st = -
read i addg to  n= abсd/e-*fgh*+%+i
All expression read
add top stack item to n   n =  abсd/e-*fgh*+%+i-
st empty

Answer:abсd/e-*fgh*+%+i-

For example a = 1 b =2 c= 64 d= 4 e = 5 f = 6 g= 7 h = 8 i = 9
a+b*(c/d-e)%(f+g*h)-i= 1+2*(64/4-5)%(6+7*8)-9= 14
abсd/e-*fgh*+%+i-
1 2 64 4/5 -*6 7 8 *+%+ 9-
stack 
4 64 2 1 
/
5 16 2 1 
-
11 2 1
*
22 1
8 7 6 22 1
*
56 6 22 1
+
62  22 1
%
22 1 
+ 
23
9 23 
- 
14